A step by step solution would really help. These are practice problems similar to my homework and I cant seem to figure them out. Thank you.
Answer #1

1)

Molar mass of O2 = 32 g/mol

mass of O2 = 46 g

mol of O2 = (mass)/(molar mass)

= 46/32

= 1.438 mol

According to balanced equation

mol of MgO formed = (2/1)* moles of O2

= (2/1)*1.438

= 2.875 mol

Molar mass of MgO,

MM = 1*MM(Mg) + 1*MM(O)

= 1*24.31 + 1*16.0

= 40.31 g/mol

mass of MgO = number of mol * molar mass

= 2.875*40.31

= 116 g

Answer: 116 g

2)

% yield = actual mass*100/theoretical mass

= 38.0*100/116

= 32.8 %

Answer: 32.8 %